OpenStudy (anonymous):

Solve . (1-cos4)*sin(2x)=cos^2(2x)

OpenStudy (anonymous):

\[(1-\cos4)*\sin(2x)=\cos^2(2x) \] \[2\cos^22x \times \sin2x-\cos^22x=0\] \[\cos^2x(2\sin2x-1)=0\]

OpenStudy (anonymous):

How come 1-cos4=2cos^2(2x) ?

OpenStudy (anonymous):

So there was no x in it thousht you missed it else convert cos^x to 1-sin^x and solve using quadratic eq